# Strong-field Gravitational Wave Lensing in the Kerr Background

### v1

@@ II.1 Incident Wave
Moreover, owing to the axisymmetry of the system, we are free to rotate about the the $z$-axis and hence, the spatial wave vector $\vec{k}$ can always be made to lie in the xzx-z plane in the initial asymptotic rest frame, at an angle $\gamma$ to the $z$-axis. In other words, $\vec{k}=\omega(\sin\gamma,0,\cos\gamma)\equiv\omega\,\hat{k}$, with $\omega>0$ being the frequency. Now, to ensure the transverse-traceless (TT) gauge, the spin-$1$ polarization vectors $\varepsilon_{\pm}^{\mu}$ must be spatial ($\varepsilon_{\pm}.\partial_{t}=0$), null ($\varepsilon_{\pm}.\varepsilon_{\pm}=0$), and transverse ($k.\varepsilon_{\pm}=0$) to the incident wave four-vector $k^{\mu}=(\omega,\vec{k})$. These conditions uniquely fix (apart from an overall normalization) the polarization vectors as $\varepsilon^{\mu}_{\pm}=(0,\cos\gamma,\pm i,-\sin\gamma)$.
(a) Parity decomposition: As the incident plane wave approaches the Kerr BH lens, it will scatter and a portion of the scattered wave will reaching the observer. The scattering problem is conveniently handled in the parity basis, rather than the helicity basis, as the parity modes do not mix with each other due to the parity-invariant [under ${\rm P}\mathrel{\mathop{\ordinarycolon}}(t,\vec{x})\to(t,-\vec{x})$] nature of the background Kerr metric. Going forward, we shall work in the parity basis. These parity modes can be calculated from Eq. (1) by applying the operators $(1\pm{\rm P})/2$ as
@@
(c) Spheroidal harmonic decomposition: The next step is to decompose $\psi_{4}^{\rm P}$ into spheroidal harmonics, which are the natural bases for Kerr perturbation. To obtain the final result in s suitable form, let us first rewrite
@@
The sum runs over $l\in[2,\infty)\cup\mathbb{Z}$ and $m\in[-l,l]\cup\mathbb{Z}$ for any fixed $l$. Now, the the asymptotic forms ($r\rightarrow\infty$) the coefficients $\widetilde{\psi}^{lm}_{\rm P}$ can be computed using the stationary phase approximation (SPA) in the limit $\omega^{\prime}r\gg 1$ as outlined in detail in Appendix-A (see also Ref. [Bautista:2022wjf]). Here, we shall only quote the final results:
@@
Note, we have replaced $r$ by the tortoise coordinate $r_{*}$ in the exponent, which is allowed as $r\to r_{*}$ in spatial infinity. This replacement is useful to identify the outgoing wave at $r\to\infty$ in its most natural form. Note also that $\psi_{4}$ also contains an ingoing piece at spatial infinity, which decays very rapidly (as $r^{-5}$) and not of our interest for the scattering problem. Hence, to make it self-evident, we shall represent the above-calculated $\psi_{4}$ by $\psi_{4}^{\rm out}$.
@@ II.3 Scattered wave at observer
$h_{\rm SW}^{\mu\nu}=\Re(A\,\xi_{+2}^{\mu}\,\xi^{\nu}_{+2}\,e^{iq\cdot x}+B\,\xi_{-2}^{\mu}\,\xi^{\nu}_{-2}\,e^{iq\cdot x}),$ (15)
@@
$\psi_{4}^{\rm SW}$ $\approx-\frac{\omega^{2}}{2}[Ae^{-i\omega(t-r_{*})}+B^{*}e^{i\omega(t-r_{*})}].$ (16)
@@ III.1 Strong-field scattering factor (SFSF)
where $\hat{n}_{\rm SO}$ is the unit spatial vector from the source to the observer and we have assumed that the polarization vectors appearing in Eq. (III.1) and Eq. (26) are approximately same as $r_{\rm SL}\ll r_{\rm SO}\sim r_{\rm LO}$. Also, as mentioned earlier, $|g|\ll|f|$ for small values of $\theta_{0}\approx\pi/6$. Thus, focusing only on the contribution of $f$ for a rough (but quite good) estimation, a suitable way to quantify the relative importance of strong-field wave scattering is by the following strong-filed scattering factor (SFSF):
@@ III.2 Mismatch analysis
A more precise assessment can be obtained by computing the mismatch between the waveform including strong-field wave-scattering effects (i.e., $h_{\rm obs}^{\mu\nu}=h_{\rm dir}^{\mu\nu}+h_{\rm SW}^{\mu\nu})$ and the corresponding waveform without these effects (i.e., $h_{\rm dir}^{\mu\nu}$). The mismatch ($\mathscr{M}$) between two waveforms $h_{1}$ and $h_{2}$ is defined as
@@
where $\mathcal{O}(h_{1},h_{2})$ is the match, defined as the normalized overlap between the waveforms maximized over any global time ($t_{0}$) and phase ($\phi_{0}$) shifts:
@@
Here, tildes denote the Fourier-transformed waveforms in the frequency domain, and $S_{n}(\nu)$ is the noise power spectral density (PSD) of the GW detector under consideration [Finn:1992xs, Cutler:1994ys, Flanagan:1997kp, Flanagan:1997sx]. In this work, we use the advanced LIGO PSD at its A+ design sensitivity [LIGOScientific:2014pky], which is planned for use in upcoming (O5) LIGO observing runs.
For illustration, we consider a GW150914-like event, consisting of two coalescing Schwarzchild BHs of masses $m_{1}=36M_{\odot}$, and $m_{2}=29M_{\odot}$. The source is at a distance of $r_{\rm SO}=475.8\,\rm{Mpc}$. Although including cosmological effects such as redshift are straightforward, we neglect them here for simplicity. The redshift $z$ of our source is $\sim 0.12$, which is quite small and can be ignored without any major effects for this demonstration. We choose the lens mass as $M=100M_{\odot}$. Moreover, the distance between the source and lens $r_{\rm SL}$ is chosen to be $100M$ or $10^{4}M_{\odot}\sim 4\times 10^{-16}\mathrm{Mpc}$, which is much smaller than the distance between the source and observer. The source-lens line is set at an angle of $\theta_{o}=30^{\circ}$ with respect to the lens-observer line. We chose this value as $30^{\circ}$ is approximately the smallest angle where the scattering analysis presented in this work may be trusted. The effect is expected to be stronger for smaller scattering angles, as seen from the plots of differential-cross section vs. scattering angle in Fig. 3. For comparison, we also present results for $\theta_{o}=60^{\circ}$ in the bottom panel.
@@
We orient our binary source such that its angular momentum points away from the lens (i.e., antiparallel to the source-lens line), this ensures that the lens experiences an approximately plane GW with purely positive helicity ($h=+2$). We further specialize to the case of on-axis ($\gamma=0$) scattering, where the spin of the lens is parallel to the line joining the source and lens. At the observer, this corresponds to an inclination (angle between the source-observer line and the binary orbital angular momentum) of $\iota=180^{\circ}-\theta_{o}=180^{\circ}-30^{\circ}=150^{\circ}$. Finally, given the axisymmetry of the setup about the z-axis (the source-lens line which is parallel to the lens spin and the source-orbital angular momentum), we choose the azimuthal angle $\phi_{o}=0^{\circ}$ without loss of generality. The setup is visually presented in Fig. 5.
@@
The observer receives the gravitational radiation as two contributions. The first is the direct component, $h^{\mu\nu}_{\rm dir}$, which is only weakly affected by the lens because of the large impact parameter and approximately follows the straight line between source and observer. Ref. [Chan:2025wgz] employs the point-lens, scalar-wave approximation to estimate lensing effects on this part. Since the point-lens approximation breaks down for large observation angles ($\theta_{o}\sim 30^{\circ}$), we do not adopt it and treat the direct wave as essentially unaffected by the lens.
The second is the scattered component, $h^{\mu\nu}_{\rm SW}$, encoding the strong-field influences of the lens. Each frequency mode acquires an additional phase $e^{i\omega(r_{\rm LO}+r_{\rm SL}-r_{\rm SO})}$. In the time domain this corresponds to a delay of approximately $c^{-1}(r_{\rm LO}+r_{\rm SL}-r_{\rm SO})\approx c^{-1}\times 2r_{\rm SL}\sin^{2}(\theta_{o}/2)\sim 0.005~\mathrm{sec}$ , much smaller than the $\sim 0.2~\mathrm{sec}$ duration of the GW150914 signal observed by LIGO. Thus the direct and scattered waves arrive nearly simultaneously and interfere, with the modulation depending on parameters such as the lens mass and spin. The observed waveform is therefore $h^{\mu\nu}_{\rm obs}=h^{\mu\nu}_{\rm dir}+h^{\mu\nu}_{\rm SW}$.
The software package PyCBC [pycbc_github] is used to compute the waveforms at the lens and at the observer. In this work, we use the IMRPhenomD [PhysRevD.93.044007, Husa:2015iqa] approximant to generate unlensed waveforms. The waveform at the lens is treated as the incident plane wave for the scattering analysis described in earlier sections, and the scattered waveform is computed in the frequency domain as outlined in Appendix D.
@@
Table 1: Mismatch between the direct component and the total waveform as the lens spin-parameter $\chi$ is varied. For a fixed scattering angle ($\theta_{o}$), spin does not seem to significantly affect the mismatch. However, increasing $\theta_{o}$ leads to a sharp drop in mismatch. The mismatch with cross polarization appears to dominate for these configurations.
@@
Fig. 6 shows the frequency-domain waveforms with and without the effect of strong-field scattering. In the top panel, $\theta_{o}=30^{\circ}$, and thus the inclination of the source for the observer is $180-30=150^{\circ}$. At this inclination, the magnitudes of $h_{+}^{\rm dir}$ and $h_{\times}^{\rm dir}$ are very close to each other (with equality holding at $180^{\circ}$). Thus, they appear almost overlapping in the plots. A similar feature is also present for the scattered wave as visible from Eq. (17) since $|g|\ll|f|$ and the waveform at the lens has purely positive helicity as mentioned before. Thus, even in the observed waveform, which includes the scattered component, this feature remains intact and we find that the curves for $h_{+}^{\rm obs}$ and $h_{\times}^{\rm obs}$ are overlapping. As a comparison, we also show results for $\theta_{o}=60^{\circ}$, where the difference in the polarization content in $\times$ basis is more clearly visible. However, the difference between the direct and observed waveforms is much smaller (the oscillations have reduced amplitude) for $\theta_{o}=60^{\circ}$. The scale of the oscillatory features in $h_{+,\times}^{\rm obs}$ is determined by the phase generated by the time delay between the two components. As $\theta_{o}$ increases, the quantity $2r_{\rm SL}\sin^{2}(\theta_{o}/2)$ grows, producing a phase that varies more rapidly with frequency. This results in correspondingly faster oscillations in the observed waveform.
Figure 6: Plots (y-axes in log-scale) illustrating the difference between the magnitude of the total observed waveform ($h_{+,\times}^{\rm obs}$) unscattered, direct component ($h_{+,\times}^{\rm dir}$) for $\theta_{o}=30^{\circ}$ (top) and $\theta_{o}=60^{\circ}$ (bottom). For $\theta_{o}=30^{\circ}$, the ++ and $\times$ polarizations contribute with nearly equal strength, causing the $\times$ curves to overlap. At $\theta_{o}=60^{\circ}$, the difference in polarization content becomes more pronounced. However, the deviation between the direct and observed waveforms is larger for $\theta_{o}=30^{\circ}$, whereas for $\theta_{o}=60^{\circ}$ the observed waveform remains close to the direct one, showing only small oscillations about it. The scale of oscillations in the observed waveform is set by the time delay between the direct and scattered components, i.e., a larger delay (corresponding to larger $\theta_{o}$) produces a greater number of oscillations.
In Table 1, we tabulate results for the mismatch in the frequency range $18,625]\,\text{Hz}$ between the plus-polarized direct component and the observed waveform, as a function of the lens-spin parameter $\chi$ for $\theta_{o}=\{30^{\circ},60^{\circ}\}$. As expected from Fig. 4 and demonstrated clearly in Fig. 6, the variation of lens spin parameter induce only very small affects in the mismatch for on-axis scattering. However, increasing the angle $\theta_{o}$ from $30^{\circ}$ to $60^{\circ}$ leads to a sharp drop in the mismatch. This is also expected from the plots in Fig. 3. This hints that it would be particularly interesting to probe $\theta_{o}\approx 0$, where the effect could be much stronger, as also suggested in Ref. [Chan:2025wgz]. Unfortunately, the formalism presented in this work for treating the scattering problem is not valid for smaller angles. Additionally, the modification of the direct component due to lensing can not be ignored for such small angles. In fact, the very notion of splitting the observed waveform into direct and scattered components would no longer be feasible. We leave the numerical study of small $\theta_{o}$ for future work. Let us also mention that a mismatch of roughly 1%, which we find for $\theta_{o}=30^{\circ}$, lies close to the detectability threshold of the current ground-based detectors and will be unambiguously detectable for future detectors. Such percentage-level mismatches are within the same order-of-magnitude as claimed in Ref. [Chan:2025wgz] for $\times$ polarization, albeit their mismatch for ++ polarization is significantly smaller. We do not find such a significant difference between the mismatches for the two polarizations.

### v2

@@ II.1 Incident Wave
Moreover, owing to the axisymmetry of the system, we are free to rotate about the the $z$-axis and hence, the spatial wave vector $\vec{k}$ can always be made to lie in the $xz$-plane in the initial asymptotic rest frame, at an angle $\gamma$ to the $z$-axis. In other words, $\vec{k}=\omega(\sin\gamma,0,\cos\gamma)\equiv\omega\,\hat{k}$, with $\omega>0$ being the frequency. Now, to ensure the transverse-traceless (TT) gauge, the spin-$1$ polarization vectors $\varepsilon_{\pm}^{\mu}$ must be spatial ($\varepsilon_{\pm}.\partial_{t}=0$), null ($\varepsilon_{\pm}.\varepsilon_{\pm}=0$), and transverse ($k.\varepsilon_{\pm}=0$) to the incident wave four-vector $k^{\mu}=(\omega,\vec{k})$. These conditions uniquely fix (apart from an overall normalization) the polarization vectors as $\varepsilon^{\mu}_{\pm}=(0,\cos\gamma,\pm i,-\sin\gamma)$.
(a) Parity decomposition: As the incident plane wave approaches the Kerr BH lens, it will scatter and a portion of the scattered wave will reach the observer. The scattering problem is conveniently handled in the parity basis, rather than the helicity basis, as the parity modes do not mix with each other due to the parity-invariant [under ${\rm P}\mathrel{\mathop{\ordinarycolon}}(t,\vec{x})\to(t,-\vec{x})$] nature of the background Kerr metric. Going forward, we shall work in the parity basis. These parity modes can be calculated from Eq. (1) by applying the operators $(1\pm{\rm P})/2$ as
@@
(c) Spheroidal harmonic decomposition: The next step is to decompose $\psi_{4}^{\rm P}$ into spheroidal harmonics, which are the natural bases for Kerr perturbation. To obtain the final result in a suitable form, let us first rewrite
@@
The sum runs over $l\in[2,\infty)\cap\mathbb{Z}$ and $m\in[-l,l]\cap\mathbb{Z}$ for any fixed $l$. Now, the the asymptotic forms ($r\rightarrow\infty$) the coefficients $\widetilde{\psi}^{lm}_{\rm P}$ can be computed using the stationary phase approximation (SPA) in the limit $\omega^{\prime}r\gg 1$ as outlined in detail in Appendix-A (see also Ref. [Bautista:2022wjf]). Here, we shall only quote the final results:
@@
Note, we have replaced $r$ by the tortoise coordinate $r_{*}$ in the exponent, which is allowed as $r\to r_{*}$ in spatial infinity. This replacement is useful to identify the outgoing wave at $r\to\infty$ in its most natural form. Note also that $\psi_{4}$ also contains an ingoing piece at spatial infinity, which decays very rapidly (as $r^{-5}$) which can be obtained from the outgoing piece using the Teukolsky-Starobinsky identities as outlined in Appendix-B. Hence, to make it self-evident, we shall represent the above-calculated $\psi_{4}$ by $\psi_{4}^{\rm out}$.
@@ II.3 Scattered wave at observer
$h_{\rm SW}^{\mu\nu}=\Re\left(A\,\xi_{+2}^{\mu}\,\xi^{\nu}_{+2}\,e^{iq\cdot x}+B\,\xi_{-2}^{\mu}\,\xi^{\nu}_{-2}\,e^{iq\cdot x}\right),$ (15)
@@
$\psi_{4}^{\rm SW}$ $\approx-\frac{\omega^{2}}{2}\left[Ae^{-i\omega(t-r_{*})}+B^{*}e^{i\omega(t-r_{*})}\right].$ (16)
@@ III.1 Strong-field scattering factor (SFSF)
where $\hat{n}_{\rm SO}$ is the unit spatial vector from the source to the observer and we have assumed that the polarization vectors appearing in Eq. (III.1) and Eq. (26) are approximately same as $r_{\rm SL}\ll r_{\rm SO}\sim r_{\rm LO}$. Also, as mentioned earlier, $|g|\ll|f|$ for small values of $\theta_{0}\approx\pi/6$. Thus, focusing only on the contribution of $f$ for a rough (but quite good) estimation, a suitable way to quantify the relative importance of strong-field wave scattering is by the following strong-field scattering factor (SFSF):
@@ III.2 Mismatch analysis
In this section, we assess the distinguishability of the lensed waveform (the waveform including strong-field wave-scattering effects) from the unlensed one (without the scattering effects), and the prospects of its detectability.
III.2.1 Setup
A natural measure of the distinguishability between two waveforms, say $h_{1}$ and $h_{2}$, is the mismatch, defined as [Dhurandhar:1992mw]:
@@ III.2.1 Setup
where $\mathcal{O}(h_{1},h_{2})$ is the match, defined as the normalized overlap between the waveforms maximized over any global time ($t_{0}$) and phase ($\phi_{0}$) shifts [Allen:2005fk, Usman:2015kfa]:
@@
Here, tildes denote the Fourier-transformed waveforms in the frequency domain, and $S_{n}(\nu)$ is the noise power spectral density (PSD) of the GW detector under consideration [Finn:1992xs, Cutler:1994ys, Flanagan:1997kp, Flanagan:1997sx]. The normalization of the waveforms is done by dividing them by their norm, $h||=\sqrt{\braket{h|h}}\equiv\rho_{\rm opt}$, where $\rho_{\rm opt}$ defines the optimal SNR.
Since in a GW search, the goal is essentially to find the template-bank member that has the smallest mismatch with the detector data, we are therefore interested in the mismatch optimized over all of the members of the unlensed waveform family. For a target waveform $h_{\rm target}$ whose match we wish to maximize using the template waveforms $h_{\rm template}$, we define fitting factor $(\textrm{FF})$ as the maximized match [Lindblom:2008cm]:
$\textrm{FF}(h_{\rm target})=\max_{\Theta_{\rm opt}}\mathcal{O}(h_{\rm template}(\Theta_{\rm opt}),h_{\rm target}),$ (32)
where, motivated by real GW searches, the maximization is done over the 44D space of binary’s intrinsic parameters ($\Theta_{\rm opt}$): chirp mass, mass ratio, and the aligned spin components. The fully optimized mismatch is, therefore, $1-\textrm{FF}$.
@@
For illustration, we consider a GW150914-like event, consisting of two coalescing Schwarzchild BHs of masses $m_{1}=36.9~\mathrm{M}_{\odot}$, and $m_{2}=32.8~\mathrm{M}_{\odot}$ [^4] We use the maximum-likelihood values of the 1D-marginalized posteriors from the ‘C01:Mixed’ channel of ‘IGWN-GWTC2p1-v2-GW150914_095045_PEDataRelease_mixed_cosmo.h5’. Posterior samples are available at Zenodo. [LIGOScientific:2016aoc, LIGOScientific:2018mvr, KAGRA:2021vkt]. The source is at a distance of $r_{\rm SO}=402.6~\rm{Mpc}$. Although including cosmological effects such as redshift are straightforward, we neglect them here for simplicity. The redshift of the source is $0.12$, which is quite small and can be ignored without any major effects for this demonstration. We choose the lens mass as $M=100~\mathrm{M}_{\odot}$. Moreover, the distance between the source and lens $r_{\rm SL}$ is chosen to be $100M$ or $10^{4}~\mathrm{M}_{\odot}\sim 4\times 10^{-16}\mathrm{Mpc}$, which is much smaller than the distance between the source and observer. We orient our binary source such that its angular momentum points away from the lens (i.e., antiparallel to the source-lens line), this ensures that the lens experiences an approximately plane GW with purely positive helicity ($h=+2$). We further specialize to the case of on-axis ($\gamma=0$) scattering, where the spin of the lens is parallel to the line joining the source and lens. The source-lens line is set at an angle of $\theta_{o}$ with respect to the lens-observer line. At the observer, this corresponds to an inclination (angle between the source-observer line and the binary orbital angular momentum) of $\iota=\pi-\theta_{o}$. Finally, given the axisymmetry of the setup about the z-axis (the source-lens line which is parallel to the lens spin and the source-orbital angular momentum), we choose the azimuthal angle $\phi_{o}=0^{\circ}$ without loss of generality. The setup is visually presented in Fig. 5.
The observer receives the gravitational radiation as two contributions. The first is the direct component, $h^{\mu\nu}_{\rm dir}$, which is only weakly affected by the lens because of the large impact parameter and approximately follows the straight line between source and observer. Ref. [Chan:2025wgz] employs the point-lens, scalar-wave approximation to estimate lensing effects on this part. Since the point-lens approximation breaks down for large observation angles (i.e., for large $\theta_{o}$, where small-angle approximations break down), we do not adopt it and treat the direct wave as essentially unaffected by the lens. The second is the scattered component, $h^{\mu\nu}_{\rm SW}$, encoding the strong-field influences of the lens. Each frequency mode acquires an additional phase $e^{i\omega(r_{\rm LO}+r_{\rm SL}-r_{\rm SO})}$. In the time domain this corresponds to a delay of approximately $c^{-1}(r_{\rm LO}+r_{\rm SL}-r_{\rm SO})\approx c^{-1}\times 2r_{\rm SL}\sin^{2}(\theta_{o}/2)\sim 0.005~\mathrm{sec}$, much smaller than the $\sim 0.2~\mathrm{sec}$ duration of the GW150914 signal observed by LIGO. Thus, the direct and scattered waves arrive nearly simultaneously and interfere, with the modulation depending on parameters such as the lens mass and spin. The observed waveform is, therefore,
$h^{\mu\nu}_{\rm obs}=h^{\mu\nu}_{\rm dir}+h^{\mu\nu}_{\rm SW}.$ (33)
For more details see Appendix-D.
We use the software package PyCBC [pycbc_github] to compute the waveforms at the lens and at the observer, and utilize the approximant IMRPhenomD [PhysRevD.93.044007, Husa:2015iqa] to generate unlensed waveforms. We use the advanced LIGO PSD at its A+ design sensitivity [LIGOScientific:2014pky], which is the target sensitivity for the upcoming (O5) LIGO observing runs. To compute the FF values, we use the Nelder-Mead algorithm [Nelder:1965zz], as implemented in the optimization module of the Scipy library [Virtanen:2019joe], to maximize the match between the waveforms. It is important to note that since we employ a maximization algorithm to compute FF, it is not sensitive to the discrete placement of templates. The waveform at the lens is treated as the incident plane wave for the scattering analysis described in earlier sections, and the scattered waveform is computed in the frequency domain as outlined in Appendix-D. We restrict ourselves to cases with $\theta_{o}>\pi/6$ as it is approximately the smallest angle where the scattering analysis presented in this work may be trusted. The effect is expected to be stronger for smaller scattering angles, as seen from the plots of differential-cross section vs. scattering angle in Fig. 3.
@@
Figure 6: Plots (y-axes in log-scale) illustrating the difference between the magnitude of the total observed waveform ($h_{+,\times}^{\rm obs}$) unscattered, direct component ($h_{+,\times}^{\rm dir}$) for $\theta_{o}=30^{\circ}$ (top) and $\theta_{o}=60^{\circ}$ (bottom). For $\theta_{o}=30^{\circ}$, both polarizations contribute with nearly equal strength, causing the $\times$ curves to overlap. At $\theta_{o}=60^{\circ}$, the difference in polarization content becomes more pronounced. However, the deviation between the direct and observed waveforms is larger for $\theta_{o}=30^{\circ}$, whereas for $\theta_{o}=60^{\circ}$ the observed waveform remains close to the direct one, showing only small oscillations about it. The scale of oscillations in the observed waveform is set by the time delay between the direct and scattered components, i.e., a larger delay produces a greater number of oscillations.
Figure 7: Variation in the minimum mismatch ($1-\textrm{FF}$), see Eq. (32), across the parameter space of $\theta_{o}$ (scattering angle), and the lens spin $\chi$, for the ++ and the $\times$ polarizations of the lensed waveform that reaches the observer after scattering shown in the left and right panels, respectively. Mismatches are computed against unlensed templates and minimised over component masses and aligned spins.
Figure 8: Comparison of the optimal SNRs of the lensed and unlensed waveforms for the ++ and $\times$ polarizations, shown in the left and right panels, respectively.
Figure 9: Illustrative estimation of the SNR required to identify scattering-induced lensing features in $h^{+}_{\rm obs}$ and $h^{\times}_{\rm obs}$, shown in the left and right panels, respectively.
@@
Table 1: Mismatch between the direct component and the total waveform as the lens spin-parameter $\chi$ is varied. For a fixed scattering angle ($\theta_{o}$), spin does not seem to significantly affect the mismatch. This is seen more clearly in the contour plots in Fig. 7. However, increasing $\theta_{o}$ leads to a sharp drop in mismatch. The mismatch with cross polarization appears to dominate for these configurations.
III.2.2 Results
@@
Fig. 6 shows the frequency-domain waveforms with and without the effect of strong-field scattering. In the top panel, $\theta_{o}=30^{\circ}$, and thus the inclination of the source for the observer is $180-30=150^{\circ}$. At this inclination, the magnitudes of $h_{+}^{\rm dir}$ and $h_{\times}^{\rm dir}$ are very close to each other (with equality holding at $180^{\circ}$). Thus, they appear almost overlapping in the plots. A similar feature is also present for the scattered wave as apparent from Eq. (17) since $|g|\ll|f|$ and the waveform at the lens has purely positive helicity as mentioned before. Thus, even in the observed waveform, which includes the scattered component, this feature remains intact and we find that the curves for $h_{+}^{\rm obs}$ and $h_{\times}^{\rm obs}$ are overlapping. As a comparison, we also show results for $\theta_{o}=60^{\circ}$, where the difference in the polarization content in $\times$ basis is more clearly visible. However, the difference between the direct and observed waveforms is much smaller (the oscillations have reduced amplitude) for $\theta_{o}=60^{\circ}$. The scale of the oscillatory features in $h_{+,\times}^{\rm obs}$ is determined by the phase generated by the time delay between the two components. As $\theta_{o}$ increases, the quantity $2r_{\rm SL}\sin^{2}(\theta_{o}/2)$ grows, producing a phase that varies more rapidly with frequency. This results in faster oscillations in the observed waveform.
In Fig. 7, we show the variation in the minimum mismatch ($1-\textrm{FF}$), see Eq. (32), across the parameter space of $\theta_{o}$ and $\chi$ for both $h^{+}_{\rm obs}$ and $h^{\times}_{\rm obs}$. As expected from Fig. 4 and demonstrated clearly in Fig. 6, the variation of lens spin parameter induce only very small affects in the mismatch for on-axis scattering. However, increasing the angle $\theta_{o}$ from $\pi/6$ to $\pi/3$ leads to a sharp drop in the mismatch. This is also expected from the plots in Fig. 3. This hints that it would be particularly interesting to probe $\theta_{o}\approx 0$, where the effect could be much stronger, as also suggested in Ref. [Chan:2025wgz]. In the case of $h^{+}_{\rm obs}$, the mismatch values decreases monotonically with increasing $\theta_{o}$, from $\sim 1\%$ at $\pi/6$ to $\sim 0.1\%$ at $\pi/2$, for all values of the lens spin as seen from Fig. 7. In contrast, the mismatch values for $h^{+}_{\rm obs}$ exhibit qualitatively different behaviour. From $\theta_{o}=\pi/6$ to $\pi/3$, the (decreasing) trend is similar to that observed for $h^{+}_{\rm obs}$. However, beyond $\theta_{o}\approx\pi/3$ the mismatch for $h^{\times}_{\rm obs}$ starts to increase and within a thin band around $\theta_{o}\approx\pi/2$, it increases sharply to reach a maximum value of $\sim 30\%$ at $\theta_{o}=\pi/2$. This large mismatch arises because near $\theta_{o}=\pi/2$ the direct cross-polarized component $h^{\times}_{\rm dir}$ vanishes [^5] This is not strictly true when higher-mode contributions are significant or when the orbital plane undergoes precession., leaving the observed signal dominated primarily by the scattered contribution, i.e., $h^{\times}_{\rm obs}=h^{\times}_{\rm dir}+h^{\times}_{\rm SW}\approx h^{\times}_{\rm SW}$.
The minimum mismatch also corresponds to the minimum fractional loss in SNR due to missing physics in the template bank [Ajith:2012mn], thereby affecting detectability. It is therefore instructive to examine how the optimal SNR itself changes in the presence of lensing. As mentioned before, optimal SNR is quantified by the norm of the total observed waveform including the scattered wave. Thus, the ratio of the optimal SNR between the lensed and unlensed waveforms captures the amplification of the waveform due to strong-field lensing. In Fig. 8, we compare the optimal SNRs of the lensed and unlensed waveforms. For the plus-polarization, we observe no significant change in the optimal SNR due to lensing, with the values remaining close to unity. In contrast, for the cross polarization, we find a substantial enhancement in the optimal SNR, up to $\sim 30\%$, near $\theta_{o}=\pi/2$, for the same reason discussed in the previous paragraph, namely the vanishing of the cross-polarization component in the direct wave for $\theta_{o}=\pi/2$.
Next, we estimate the SNR required to identify lensing features in the signal. We use the approximation for the Bayes factor in terms of the fitting factor [Cornish:2011ys, DelPozzo:2014cla],
$\ln\mathscr{B}\approx\frac{\rho^{2}}{2}\left(1-\rm{FF}^{2}\right),$ (34)
and require the lensed-versus-unlensed ln-Bayes factor to be at least $10$ for the identification of lensing features. We adopt a higher threshold than that suggested by, for example, Jeffreys’ criterion [Jeffreys:1939xee], since this approximation is known to overestimate the Bayes factor (see, e.g., Refs. [DelPozzo:2014cla, Mishra:2023ddt]). Fig. 9 shows the estimated SNR required, $\rho_{\rm req}$, to identify scattering-induced lensing features in $h^{+}_{\rm obs}$ and $h^{\times}_{\rm obs}$. We find that $\rho_{\rm req}\gtrsim 25$ near $\theta_{o}=\pi/6$, increasing to $\sim 100$ at $\theta_{o}=\pi/2$ for $h^{+}_{\rm obs}$, while it decreases to $\sim 10$ for $h^{\times}_{\rm obs}$. This behaviour follows directly from the corresponding decrease in the fitting factor values (see Fig. 7). We emphasize that the signal observed by a detector is a projection of the plus and cross polarizations. Consequently, near $\theta_{o}=\pi/2$ although scattering-induced features in $h^{\times}_{\rm obs}$ become detectable in isolation, the observed signal may still be dominated by $h^{+}_{\rm obs}$, potentially hindering their detectability. Additionally, note that although the required SNR is lower for $\theta_{o}=\pi/2$ when the cross-polarization is considered in isolation, detecting a signal that meets this SNR threshold may in fact be more difficult. This is because $h^{\times}_{\rm obs}$ is dominated by the scattered waveform for $\theta_{o}=\pi/2$, which has a smaller amplitude than the unlensed waveform due to the increased travel time and the corresponding additional decay [see Eq. (28)], as well as the directional dependence of the scattering amplitude, see Fig. 3. Nevertheless, with future detectors it may be possible to observe signals in this regime with SNRs exceeding the required threshold.
Finally, we note that the scattering formalism presented in this work is not valid for small angles ($\theta_{o}\lesssim\pi/6$), where the modification of the direct component due to lensing cannot be ignored. In fact, the very notion of splitting the observed waveform into direct and scattered components would no longer be feasible. Nevertheless, the convergence of the partial-wave expansion might be improved by adopting an eikonal approach. Such an improvement has indeed been demonstrated recently [CarrilloGonzalez:2025gqm] by approximating the sum over $l$-modes by an integral over the impact parameter. We leave the numerical study of small $\theta_{o}$ for future work.